What positive number added to its reciprocal gives the minimum sum?


1
Expert's answer
2022-03-25T06:28:01-0400

"x+\\frac{1}{x}\\rightarrow \\min \\\\Cauchy\u0091s\\,\\,inequality:\\\\x+\\frac{1}{x}\\geqslant 2\\sqrt{x\\cdot \\frac{1}{x}}=2\\\\The\\,\\,equality\\,holds\\,for\\,\\,x=\\frac{1}{x}\\Rightarrow x=1\\\\Answer: 1"